Complete the factorization
16x^7 = 2x^2( )
32x^5
8x^4
8x^5
32^x
this is multiple choice but can you work it out for me so i can get it?

Answer by checkley77(12844)

16x^7 = 2x^2(Z)
Z=16x^7/2x^2
Z=8x^5
16x^7=2x^2(8x^5)
16x^7=16x^7
32x^5
8x^4
8x^5 answer.
32^x